OpenStudy (anonymous):

find f divided by g f(x) = sqrt (x-5) and g(x) sqrt (x)

OpenStudy (bahrom7893):

no it's: sqrt((x-25)/x)

OpenStudy (anonymous):

y?

OpenStudy (bahrom7893):

f(x)=sqrt(x-5) g(x)=sqrt(x) f(x)/g(x)=sqrt(x-5)/sqrt(x)=sqrt((x-25)/x)

OpenStudy (anonymous):

and if u square that?

OpenStudy (bahrom7893):

he just said f(x)/g(x)

OpenStudy (bahrom7893):

ur answer is: (f(x)/g(x))^2

OpenStudy (anonymous):

respect

OpenStudy (anonymous):

thanks alot!! so the x on the bottom isn't sqrtx its just x?

OpenStudy (bahrom7893):

yea.. but the square root is over the whole thing

OpenStudy (bahrom7893):